OpenStudy (anonymous):

Evaluate the definite integral: int_(0)^(3) 6(t^5 - t^3 + 1) \ dt

OpenStudy (anonymous):

\[ \int_0^3 6(t^5 - t^3 + 1) dt =\left[ 6 \left( \frac {t^6}6- \frac {t^4}4 +t \right) \right]_0^3 \]

OpenStudy (anonymous):

You should be able to finish it.